What is the standard form for this conic section?
X^2+16y^2+12x+160y+276=0

Answer by Alan3354(69443)

What is the standard form for this conic section?
X^2+16y^2+12x+160y+276=0
=============
Complete the squares for x and for y separately.
X^2+16y^2+12x+160y+276=0
X^2+12x + 16y^2+160y = -276